32. In Japan, taxes and real imports do not depend on real income. Autonomous real consumption is ¥400 million, lump-sum taxeIn Japan, taxes and real imports do not depend on real income. Autonomous real consumption is ¥400 million, lump-sum taxes (taxes that do not depend on real income) are ¥100 million, investment spending is ¥300 million, Japanese government spending is ¥100, and real net exports are ¥0, The Japanese Marginal Propensity to Consume is 0.60. a. Solve for the equilibrium level of Japanese real GDP. b. Suppose that more foreign investors begin to buy Japanese stocks on the Tokyo Stock Exchange, and the NIKKEI 225 Stock Market Index (the Japanese Dow-Jones index) increases. by 600 points. what do you expect to happen to Japanese real consumption? c. If Japanese autonomous real consumption increases by ¥150 million, what is the new level of equilibrium Japanese real GDP?
Answer #1

Ans:a) Given, consumption spending= 400 million

Investment spending= 300 million

Government spending=100 million

Real net exports= 0, MPC= 0.60

Equilibrium GDP = Consumpumtion spending+ Investment spending + Governement spending + Net exports

=400+300+100= 800 Million

Ans:b) This effect wil indrectly increase the consumption spending and direct increase the investment spending.

Ans:c) The equilibrium GDP will increase by 150 million, i.e 800+150= 950 Million